# FlexGuard: A Design Space for On-Body Feedback for Injury Prevention in Strength Training

### v1

@@ Abstract.
Strength training carries risk of injury when exercises are performed without supervision. While haptics research has advanced, there remains a gap in how to integrate on-body feedback into intelligent wearables. Developing such a design space requires experiencing feedback in context, yet obtaining functional systems is costly. By addressing these challenges, we introduce FlexGuard, a design space for on-body feedback to support injury prevention in strength training. The design space was derived from nine co-design workshops, where novice trainees and expert trainers DIY’d low-fidelity on-body feedback systems, tried them immediately, and surfaced needs and challenges encountered in real exercising contexts. We then evaluated the space through speed dating, using storyboards to cover the design dimensions. We followed up with workshops to further validate selected dimensions in practice through a proof-of-concept wearable system prototype. Our findings extend the design space for sports and fitness wearables in the context of strength training.
@@ 1. Introduction
Strength training is a widely practiced method for improving strength, mobility, and overall health. However, errors such as joint misalignment or overextension can increase the risk of injury, especially for novice practitioners (Fisher, 2025). Personal trainers help mitigate these risks by providing tailored feedback, but such supervision is costly and depends on others’ availability, making it difficult to scale. Advances in haptics now enable the emulation of sensations such as taps, pressure, and resistance. This creates an opportunity for intelligent wearable systems that can deliver embodied feedback to support safe and effective training. What remains missing, however, is the design knowledge needed to integrate on-body feedback into such systems.
@@
Designing on-body feedback systems is challenging because users have to physically experience feedback during strength training (Tennent et al., 2020), and obtaining functional systems is costly (Caserman et al., 2021). Investing heavily in a prototype to serve as a technology probe (Hutchinson et al., 2003) is often impractical in early-stage exploration. Moreover, any probing system must be representative of the diverse on-body feedback mechanisms that have been studied and developed. Together, these challenges hinder systematic investigation of the design space.
@@
By addressing these challenges, we developed FlexGuard, a design space for on-body feedback in strength training injury prevention. We began by surveying the landscape of feedback modalities, including vibrotactile (Huang et al., 2025), pneumatic (Delazio et al., 2018; Pardomuan et al., 2024), and other approaches (Kwon et al., 2022; Lin et al., 2025b), and grouped them into two categories: extrinsic, which require active actuation (e.g., vibration, air pressure), and intrinsic, which rely on material properties (e.g., orthoses, braces) to provide passive feedback once configured. We then prepared technology probes representing these categories, along with supporting props, and conducted nine co-design workshops in our institution’s gym. Each session paired a novice trainee with an expert trainer, who collaboratively DIY’d personalized low-fidelity on-body feedback systems for five representative exercises and tested them immediately in practice. Through thematic analysis of session recordings and worksheets, we distilled the needs and challenges that informed the derivation of FlexGuard.
@@
FlexGuard is organized around two complementary dimensions: Sensing & Triggering and Feedback Delivery. Sensing & Triggering specifies which indicators should be monitored (e.g., joint angle, movement dynamics, motion path, muscle activation, inter-limb coordination), how triggering policies are defined (e.g., expert-based or personalized), and where monitoring should occur (e.g., target, stabilizing, moving, or compensation regions). Feedback Delivery describes how feedback is provided once triggered, including the body zones where different feedback types are applied, the intensities of feedback, its directional strategies, and its timing across movement phases. Together, these dimensions define how on-body feedback can be timed, targeted, and configured.
We ran three speed dating workshops with seven trainees with limited strength training experience. Participants first performed exercises with the provided equipment in our lab to ground their reflections, and then discussed a series of storyboards that varied Sensing & Triggering and Feedback Delivery, spanning the dimensions of FlexGuard. The results revealed that participants interpreted feedback modalities with distinct semantics, emphasized the importance of feedback timing and intensity, and highlighted adaptability as critical across phases of movement, exercises, and user states. They also reflected on learning and safety, noting concerns that excessive support could reduce learning and that locking mechanisms could pose risks. Overall, these findings validated the proposed dimensions of FlexGuard while also pointing to areas for refinement.
@@
We further validated the dimensions of FlexGuard in practice by developing a proof-of-concept (PoC) wearable system. Motivated by the speed dating findings, the PoC focused specifically on adaptability. Given the underexplored nature of adaptive intrinsic feedback, we narrowed the design to intrinsic mechanisms, implementing a strap-based suit tested on two exercises that together cover abduction–adduction and flexion–extension degrees of freedom of the shoulder. A camera-based system detected deviations from target posture and signaled strap adjustments, which were manually enacted in this PoC to simulate an automated system. This setup allowed participants to experience phase-specific variations in feedback direction and intensity, providing finer-grained validation of how the design space principles translate into embodied training contexts at the phase level.
@@
We introduce FlexGuard, a design space for on-body feedback in strength training, grounded in nine co-design workshops with trainers and trainees.
We articulate dimensions that characterize how systems can sense and deliver feedback to reduce injury risk across the exercises studied in the workshops.
@@ 2. Related Work
Our research builds upon prior work in four key areas: Injury Prevention in Strength Training, Sensing and Feedback Systems for Physical Training, Advances in Materials and Wearable Technologies, Intelligent Systems in Movement and Skill Learning.
2.1. Injury Prevention in Strength Training
@@
Strength training carries inherent risks of injury, particularly when exercises involve high external loads that stress the musculoskeletal system. Body regions that both bear the load and move dynamically, such as the lower back and shoulders, are especially prone to injuries like strains, sprains, and impingements (Hedrick and Wada, 2008; Barnes et al., 2021). While such injuries are common, they are often avoidable. Research indicates that a large portion of lower back injuries are linked to improper lifting form, excessive loads, and overtraining—all factors that can be mitigated through proper technique and training practices (Alqarni, 2019). Faulty alignment and compensatory movement patterns introduce abnormal stress across joints and soft tissues, contributing not only to acute injury but also to long-term issues. This risk is especially pronounced in novice trainees, whose motor control systems may not yet be developed enough to coordinate complex compound movements effectively (Aslam et al., 2025). Thus, consistent attention to motor learning is fundamental for safe and effective strength training (Latella et al., 2019).
@@
This is why working with a personal trainer is widely recommended, particularly for beginners. Professional supervision ensures that trainees are taught correct movement techniques, are properly spotted to maximize repetitions safely, and are guided through appropriate loading progressions. When properly coached, strength training is a safe activity with a low incidence of injury relative to other sports (Fisher et al., 2023). In contrast, unsupervised training is associated with higher risks due to inconsistent technique and overly aggressive load selection (Faigenbaum and Myer, 2010; Gabbett, 2020). While technologies exist to support self-coaching (Snouzy, 2025), human coaches provide multifaceted guidance that extends beyond form correction and motor habit development to include enforcing safety practices such as spotting, load monitoring, and early detection of execution breakdowns (Fisher et al., 2023).
@@
In addition to coaching, trainees frequently use supportive wearable gear to enhance injury prevention. Equipment such as weightlifting belts, wrist wraps, knee sleeves, and elbow sleeves is commonly used in gyms to help maintain alignment and distribute stress during heavy lifts. Weightlifting belts, for instance, are used to increase intra-abdominal pressure and reduce spinal loading during squats and deadlifts (Herbaut and Tuloup, 2025). Knee sleeves improve proprioception during deep knee flexion movements (Machek et al., 2021). Users can also define and produce custom wearables for their specific needs (Xu et al., 2025). For example, tailoring their own knee patch or squat suit to support heavy lifting (Latella et al., 2024).
@@
FlexGuard builds on this trajectory by extending injury prevention beyond passive support and human supervision. Leveraging haptics for on-body sensation, FlexGuard advances toward adaptive systems that actively sense and respond to cues and deviations during exercise, providing intuitive feedback to promote safer training.
@@ 2.2. Sensing and Feedback Systems for Physical Training
Designing wearables requires careful consideration of how the body perceives and responds to interaction, emphasizing not just function but felt experience (Tennent et al., 2020); thus, there are two key components: sensing and feedback methods on the body. Sensing technologies have enabled researchers and designers to monitor a wide range of body states during physical training. Researchers have explored the use of sensors to monitor muscle activity (Zhu et al., 2022; Esposito et al., 2018; Rudolph et al., 2022). Systems like RecoFit (Morris et al., 2014) and LiftSmart (Kowsar et al., 2019) use wearable sensors to monitor and assess users’ exercise performance. Other systems rely on pressure or tactile sensors to provide posture-related sensing, such as GymSoles, which uses insole-based pressure readings to visualize a user’s center of pressure during lifts (Elvitigala et al., 2019). FormFit (Rao et al., 2024) employs computer vision for automatic pose correction, while Zishi (Wang et al., 2016) combines smart textiles and motion sensing to monitor posture during rehabilitation. More integrated systems like E-Orthosis (Yang et al., 2023) embed sensing directly into orthotic devices using textile electrodes, enabling real-time feedback and data logging in rehabilitation and sports contexts. Researchers have also studied textile-based knee strain sensors for real-time monitoring during workouts (Aliyana et al., 2024), as well as developable shoulder sensors for precise movement tracking (Evke et al., 2019). Commercial platforms such as Catapult (Catapult, 2025) offer elite-level sports analytics, although they often remain inaccessible or inflexible for everyday users.
@@
A parallel stream of work focuses on feedback modalities for physical training. Researchers have extensively explored visual augmentation through projection and AR/VR (Yu et al., 2024), including systems such as BodyLights (Turmo Vidal et al., 2020), augmented instructional videos (Semeraro and Turmo Vidal, 2022), real-time overlays (Keelawat and Suzuki, 2024; Ma et al., 2024), and spatial visualizations in FlowAR (Jo et al., 2023a), all of which guide users with contextual cues. Some systems, like Weight-Mate (Paay et al., 2020), combine visual and auditory feedback to help users maintain alignment. Haptic feedback systems have also gained attention, including vibration-based alerts for physical activities (Spelmezan et al., 2009; Kosmalla et al., 2016), implemented in TrainerTap (Jo et al., 2023b) VibraForge (Huang et al., 2025) and Subtletee (Woźniak et al., 2020). Such systems may be particularly beneficial for individuals with visual impairments when practicing yoga (Islam et al., 2022). Han et al. (Han et al., 2023) explore 3D-printed tactor geometries for generating diverse haptic sensations. EMS-based stimulation for movement correction is another promising direction; for instance, Teslasuit (Caserman et al., 2021) exemplifies a production-grade electro muscle stimulation (EMS) and transcutaneous electrical nerve stimulation (TENS) platform capable of training full-body motion. Other systems explore novel modalities, including skin-stretching for proprioceptive illusions (Lin et al., 2024b), pneumatic water-jet feedback (Zhang et al., 2024), localized pressure for knee joint stabilization (Park et al., 2020), and torque-based force guidance for arm motion (Kim and Asbeck, 2022).
@@ 2.3. Advances in Materials and Wearable Technologies
Recent developments in soft robotics, programmable materials, and digital fabrication have expanded the design space for wearable systems. Pneumatic actuation systems like PneumAct (Günther et al., 2019), VabricBeads (Pardomuan et al., 2024), Force Jacket (Delazio et al., 2018), and Frozen Suit (Al Maimani and Roudaut, 2017) use air or liquid pressure to generate movement, resistance, or stiffness changes in the body. milliMorph (Lu et al., 2019) pushes this further by introducing millimeter-scale fluidic chambers for fine-grained, fluid-driven shape change. PneuMa (Saini et al., 2024) introduces pneumatic bodily extensions to support daily movement. Actuation mechanisms inspired by biological forms, such as artificial muscles (Li et al., 2017), offer compact and powerful actuation possibilities. Researchers have also investigated collaborative approaches for designing wearables with on-body feedback, such as vibrotactile patterns in remote scenarios where designers are not co-located (Wittchen et al., 2025).
Programmable materials with bistable or multistable properties allow for richer forms of mechanical feedback and passive guidance (Alu et al., 2025). Works such as Kirigami Haptic Swatches (Chang et al., 2020), FlexTure (Johnson et al., 2023), Waxpaper Actuator (Wu et al., 2024), and ConeAct (Lin et al., 2024a) explore structural snapping transitions, geometric locking, and reactive haptics. Snap-through mechanisms, like those in SnapInflatables (Yang et al., 2024), enable rapid shape changes with localized feedback. Springlets (Hamdan et al., 2019) introduces shape-memory alloy springs that simulate on-skin sensations such as touching, pinching, pulling, etc.
Wearable fabric-based sensing and actuation is also advancing rapidly (Li and Luo, 2024). Luo et al. demonstrate textile-based control for skill transfer tasks like piano playing (Luo et al., 2024), while ReKnit-Care (Hu et al., 2025) combines knitted sensors with EMS to provide fine-grained hand feedback for rehabilitation. Lin et al. (Lin et al., 2025a) introduce passive microstructures to switch between states while worn to change perceived physical environment. Tessmer et al. designed a spacesuit with tunable compression (Tessmer et al., 2022). FlexGuard draws from this material ecosystem, integrating advances in actuation and textiles to explore how adaptive, on-body feedback can support safe and personalized strength training.

### v2

# FlexGuard: A Design Space for On-Body Feedback for Safety Scaffolding in Strength Training

@@ Abstract.
Strength training carries inherent safety risks when exercises are performed without supervision. While haptics research has advanced, there remains a gap in how to integrate on-body feedback into intelligent wearables. Developing such a design space requires experiencing feedback in context, yet obtaining functional systems is costly. By addressing these challenges, we introduce FlexGuard, a design space for on-body feedback that scaffolds safety during strength training. The design space was derived from nine co-design workshops, where novice trainees and expert trainers DIY’d low-fidelity on-body feedback systems, tried them immediately, and surfaced needs and challenges encountered in real exercising contexts. We then evaluated the design space through speed dating, using storyboards to cover the design dimensions. We followed up with workshops to further validate selected dimensions in practice through a proof-of-concept wearable system prototype, examining how on-body feedback scaffolds safety during exercise. Our findings extend the design space for sports and fitness wearables in the context of strength training.
@@ 1. INTRODUCTION
Strength training is a widely practiced method for improving strength, mobility, and overall health. However, errors such as joint misalignment or overextension can increase the risk of injury, especially for novice practitioners (Fisher, 2025). Personal trainers help mitigate these risks by providing tailored feedback that cultivates trainees’ awareness of potentially unsafe movement and offers moment-to-moment physical support during execution. Such in-person supervision, however, is costly and depends on others’ availability, making it difficult to scale. As a result, there is a need for alternative forms of support that can scaffold trainees’ experience by helping them remain aware of safety-relevant movement while also providing embodied assistance during training. Advances in haptics now enable the emulation of physical sensations commonly used in in-person coaching, such as taps, pressure, and resistance. This creates an opportunity for intelligent wearable systems that can deliver on-body feedback to scaffold safety during strength training. What remains missing, however, is the design knowledge needed to integrate such on-body feedback into wearable systems.
@@
Designing on-body feedback systems is challenging because users must physically experience feedback during strength training (Tennent et al., 2020; Ley-Flores et al., 2024), and obtaining functional systems is costly (Caserman et al., 2021). Investing heavily in a prototype to serve as a technology probe (Hutchinson et al., 2003) is often impractical in early-stage exploration. Moreover, any probing system must be representative of the diverse on-body feedback mechanisms that have been studied and developed. Together, these challenges hinder systematic investigation of the design space.
@@
By addressing these challenges, we developed FlexGuard, a design space for on-body feedback to scaffold safety in strength training. We began by surveying the landscape of feedback modalities, including vibrotactile (Huang et al., 2025; Elsayed et al., 2020), pneumatic (Delazio et al., 2018; Pardomuan et al., 2024), and other approaches (Kwon et al., 2022; Lin et al., 2025b), and grouped them into two categories: extrinsic, which require active actuation (e.g., vibration or air pressure), and intrinsic, which rely on material properties (e.g., orthoses or braces) to provide passive feedback once configured. We then prepared technology probes representing these categories, along with supporting props, and conducted nine co-design workshops in our institution’s gym. Each session paired a novice trainee with an expert trainer, who collaboratively DIY’d personalized low-fidelity on-body feedback systems for five representative exercises and tested them immediately in practice. Through thematic analysis, we distilled the needs and challenges that informed the derivation of FlexGuard.
@@
FlexGuard is organized around two complementary dimensions: Sensing & Triggering and Feedback Delivery. Sensing & Triggering specifies which indicators should be monitored (e.g., joint angle, movement dynamics, motion path, muscle activation, or inter-limb coordination), how triggering policies are defined (e.g., expert-based or personalized), and where monitoring should occur (e.g., target, stabilizing, moving, or compensation regions). Feedback Delivery describes how feedback is provided once triggered, including the body zones where different feedback types are applied, the intensity of feedback, its directional strategies, and its timing across movement phases. Together, these dimensions define how on-body feedback can be timed, targeted, and configured to scaffold safety during exercise.
We ran three speed dating workshops with seven trainees with limited strength training experience. Participants first performed exercises with the provided equipment in our lab to ground their reflections and then discussed a series of storyboard scenarios that varied Sensing & Triggering and Feedback Delivery, spanning the dimensions of FlexGuard. The results indicated that participants found the speculative scenarios relevant and useful. These findings provided conceptual validation of FlexGuard and allowed us to refine the design space based on evidence of user alignment with its dimensions. n addition, we identified underlying needs across scenarios, including expectations around feedback semantics, timing and progressivity, and adaptability across users, exercises, and contexts.
@@
We further validated selected dimensions of FlexGuard in practice by developing a proof-of-concept (PoC) wearable system. Motivated by the speed dating findings, we narrowed the scope of the PoC to focus on intrinsic feedback and the underlying user needs identified earlier. As wearable systems move toward personalized intelligent support, intrinsic feedback becomes particularly relevant due to its potential for continuous and immediate embodied assistance. However, prior work provides limited design knowledge on how intrinsic mechanisms should be integrated with real movement contexts. To address this gap, we implemented a strap-based intrinsic-feedback suit tested on two exercises that together span two primary degrees of freedom of the shoulder. A camera-based system detected deviations from target posture and signaled strap adjustments, which were manually enacted in this PoC to simulate an automated system. This setup allowed participants to experience phase-specific variations in feedback direction and intensity, providing embodied validation of the design space and confirming key assumptions from the speed dating study in practice.
@@
We introduce FlexGuard, a design space for on-body feedback that scaffolds safety in strength training, grounded in nine co-design workshops with trainers and trainees.
We articulate dimensions that characterize how systems can sense and deliver feedback to scaffold trainees’ awareness of and support for safety-relevant movement during exercise.
@@ 2. RELATED WORK
Our research builds upon prior work in four key areas: Safety and Risk in Strength Training, Sensing and Feedback Systems for Physical Training, Advances in Materials and Wearable Technologies, Intelligent Systems in Movement and Skill Learning.
2.1. Safety and Risk in Strength Training
@@
Strength training carries inherent safety considerations, particularly when exercises involve high external loads that stress the musculoskeletal system. Body regions that both bear the load and move dynamically, such as the lower back and shoulders, are especially prone to overload or misalignment-related issues (Hedrick and Wada, 2008; Barnes et al., 2021). While such issues are common, they are often avoidable. Research indicates that a large portion of lower back strain incidents are linked to improper lifting form, excessive loads, and overtraining, which all are factors that can be mitigated through proper technique and training practices (Alqarni, 2019). Faulty alignment and compensatory movement patterns introduce abnormal stress across joints and soft tissues, contributing not only to unsafe strain but also to long-term issues. This risk is especially pronounced in novice trainees, whose motor control systems may not yet be developed enough to coordinate complex compound movements effectively (Aslam et al., 2025). Thus, consistent attention to motor learning is fundamental for safe and effective strength training (Latella et al., 2019).
@@
This is why working with a personal trainer is widely recommended, particularly for beginners. Professional supervision ensures that trainees are taught correct movement techniques, are properly spotted to maximize repetitions safely, and are guided through appropriate loading progressions. When properly coached, strength training is a safe activity relative to other sports (Fisher et al., 2023). In contrast, unsupervised training is associated with higher risks due to inconsistent technique and overly aggressive load selection (Faigenbaum and Myer, 2010; Gabbett, 2020). While technologies exist to support self-coaching (Snouzy, 2025), human coaches provide multifaceted guidance that extends beyond form correction and motor habit development to include enforcing safety practices such as spotting, load monitoring, and early detection of execution breakdowns (Fisher et al., 2023).
@@
In addition to coaching, trainees frequently use supportive wearable gear to enhance safe training practices. Equipment such as weightlifting belts, wrist wraps, knee sleeves, and elbow sleeves is commonly used in gyms to help maintain alignment and distribute stress during heavy lifts. Weightlifting belts, for instance, are used to increase intra-abdominal pressure and reduce spinal loading during squats and deadlifts (Herbaut and Tuloup, 2025). Knee sleeves improve proprioception during deep knee flexion movements (Machek et al., 2021). Users can also define and produce custom wearables for their specific needs (Xu et al., 2025). For example, tailoring their own knee patch or squat suit to support heavy lifting (Latella et al., 2024).
@@
FlexGuard builds on this trajectory by extending safety considerations beyond passive support and human supervision. Leveraging haptics for on-body sensation, FlexGuard advances toward adaptive systems that actively sense and respond to cues and deviations during exercise, providing intuitive feedback to support safety scaffolding during training.
@@ 2.2. Sensing and Feedback Systems for Physical Training
Designing wearables requires careful consideration of how the body perceives and responds to interaction, emphasizing not just function but felt experience (Tennent et al., 2020; Windlin et al., 2022); thus, there are two key components: sensing and feedback methods on the body. This aligns with the principles of soma-based design theory, which foregrounds the subjective, lived body as a central resource for interaction design (Höök et al., 2017). Ley-Flores et al. used a ‘somatic dress-up’ approach to explore how negative body perceptions can hinder engagement in physical activity (Ley-Flores et al., 2024). Sensing technologies have enabled researchers and designers to monitor a wide range of body states during physical training. Researchers have explored the use of sensors to monitor muscle activity (Zhu et al., 2022; Esposito et al., 2018; Rudolph et al., 2022). Systems like RecoFit (Morris et al., 2014) and LiftSmart (Kowsar et al., 2019) use wearable sensors to monitor and assess users’ exercise performance. ProxiFit (Kim et al., 2023) senses exercise repetitions by detecting magnetic variations during movement. Other systems rely on pressure or tactile sensors to provide posture-related sensing, such as GymSoles, which uses insole-based pressure readings to visualize a user’s center of pressure during lifts (Elvitigala et al., 2019). FormFit (Rao et al., 2024) employs computer vision for automatic pose correction, while Zishi (Wang et al., 2016) combines smart textiles and motion sensing to monitor posture during rehabilitation. More integrated systems like E-Orthosis (Yang et al., 2023) embed sensing directly into orthotic devices using textile electrodes, enabling real-time feedback and data logging in rehabilitation and sports contexts. Researchers have also studied textile-based knee strain sensors for real-time monitoring during workouts (Aliyana et al., 2024), as well as developable shoulder sensors for precise movement tracking (Evke et al., 2019). Commercial platforms such as Catapult (Catapult, 2025) offer elite-level sports analytics, although they often remain inaccessible or inflexible for everyday users.
@@
A parallel stream of work focuses on feedback modalities for physical training. Researchers have explored output devices such as smart rings (Wang et al., 2025), along with visual augmentation via projection and AR/VR (Yu et al., 2024). Examples include systems like BodyLights (Turmo Vidal et al., 2020), augmented instructional videos (Semeraro and Turmo Vidal, 2022), real-time overlays (Keelawat and Suzuki, 2024; Ma et al., 2024), and spatial visualizations in FlowAR (Jo et al., 2023a), all of which guide users through contextual cues during training. Some systems, like Weight-Mate (Paay et al., 2020), combine visual and auditory feedback to help users maintain alignment. Haptic feedback systems have also gained attention, including vibration-based alerts for physical activities (Spelmezan et al., 2009; Kosmalla et al., 2016), implemented in TrainerTap (Jo et al., 2023b), VibraForge (Huang et al., 2025) and Subtletee (Woźniak et al., 2020). Such systems may be particularly beneficial for individuals with visual impairments when practicing yoga (Islam et al., 2022). Han et al. (Han et al., 2023) explore 3D-printed tactor geometries for generating diverse haptic sensations. EMS-based stimulation for movement correction is another promising direction. For instance, FootStriker helps improve running foot form to prevent knee-related injuries (Hassan et al., 2017). Additionally, the Teslasuit (Caserman et al., 2021) exemplifies a production-grade EMS and TENS platform capable of training full-body motion. Wearable robotics such as HapticSnakes, which use waist-worn robotic arms to deliver multi-haptic feedback (including taps and brushing) to various locations across the upper and lower body (Al-Sada et al., 2020). Other systems explore novel modalities, including skin-stretching for proprioceptive illusions (Lin et al., 2024b), pneumatic water-jet feedback (Zhang et al., 2024), localized pressure for knee joint stabilization (Park et al., 2020), and torque-based force guidance for arm motion (Kim and Asbeck, 2022). While significant work has focused on the upper body, some systems involve skin-stretching on the legs to provide proprioceptive feedback for the lower body, demonstrating the applicability of on-body cues to other regions (Wang et al., 2020).
@@ 2.3. Advances in Materials and Wearable Technologies
Recent developments in soft robotics, programmable materials, and digital fabrication have expanded the design space for wearable systems. Pneumatic actuation systems like PneumAct (Günther et al., 2019), VabricBeads (Pardomuan et al., 2024), Force Jacket (Delazio et al., 2018), and Frozen Suit (Al Maimani and Roudaut, 2017) use air or liquid pressure to generate movement, resistance, or stiffness changes in the body. milliMorph (Lu et al., 2019) pushes this further by introducing millimeter-scale fluidic chambers for fine-grained, fluid-driven shape change. PneuMa (Saini et al., 2024) introduces pneumatic bodily extensions to support daily movement. Actuation mechanisms inspired by biological forms, such as artificial muscles (Li et al., 2017), offer compact and powerful actuation possibilities. Researchers have also investigated collaborative approaches for designing wearables with on-body feedback such as vibrotactile patterns (Wittchen et al., 2025; Sung et al., 2024; Zhou et al., 2025). Furthermore, to enhance further capability, tools like PneuMod enable computational design and fabrication of pneumatic actuators with programmable material properties, allowing for fine-tuned force and stiffness responses (Zhang and Sra, 2021).
Programmable materials with bistable or multistable properties allow for richer forms of mechanical feedback and passive guidance (Alu et al., 2025; Lee et al., 2026). Works such as Kirigami Haptic Swatches (Chang et al., 2020), FlexTure (Johnson et al., 2023), Waxpaper Actuator (Wu et al., 2024), and ConeAct (Lin et al., 2024a) explore structural snapping transitions, geometric locking, and reactive haptics. Snap-through mechanisms, like those in SnapInflatables (Yang et al., 2024), enable rapid shape changes with localized feedback. Springlets (Hamdan et al., 2019) introduces shape-memory alloy springs that simulate on-skin sensations such as touching, pinching, pulling, etc.
Wearable fabric-based sensing and actuation is also advancing rapidly (Li and Luo, 2024). SeamFit introduces garments capable of logging exercise movements (Yu et al., 2025). Luo et al. demonstrate textile-based control for skill transfer tasks like piano playing (Luo et al., 2024), while ReKnit-Care (Hu et al., 2025) combines knitted sensors with EMS to provide fine-grained hand feedback for rehabilitation. Lin et al. (Lin et al., 2025a) introduce passive microstructures to switch between states while worn to change perceived physical environment. Tessmer et al. designed a spacesuit with tunable compression (Tessmer et al., 2022). FlexGuard draws from this material ecosystem, integrating advances in actuation and textiles to explore how adaptive, on-body feedback can support safety scaffolding and personalization during strength training.